40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 235 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  5254   Wed Aug 17 12:14:27 2011 Josh SmithOmnistructureComputer Scripts / Programs40m summary page plans

Josh Smith, Fabian Magana-Sandoval, Jackie Lee (Fullerton)

Thanks to Jamie and Jenne for the tour and the input on the pages.

We had a look at the GEO summary pages and thought about how best to make a 40m summary page that would eventually become and aligo summary page. Here's a rough plan:

- First we'll check that we can access the 40m NDS2 server to get data from the 40m lab in Fullerton.

- We'll make a first draft of a 40m summary page in python, using pynds, and base the layout on the current geo summary pages.

- When this takes shape we'll iterate with Jamie, Jenne, Rana to get more ideas for measurements, layout.

Other suggestions: Jenne is working on an automated noisebudget and suggests having a placeholder for it on the page. We can also incorporate some of the features of Aidan's 40m overview medm screen that's in progress, possibly with different plots corresponding to different parts of the drawing, etc. Jenne also will email us the link of once per hour medm screenshots.

 

  5253   Wed Aug 17 06:42:38 2011 kiwamuUpdateGeneralin-vac work : the end is near

We will pump down the chambers on Thursday morning.

Today will be a day of the OSEM and oplev party.

 

 -- to do list for today --

 + OSEM mid-range adjustment

 + oplev realignment

 + placement of beam traps

 + extraction of IPPOS

 + table leveling

 + interferometer alignment

 + AM-PM mystery

 + preparation for drag and wipe

  5252   Wed Aug 17 03:10:06 2011 kiwamuUpdateGeneralcurrent status of in-vac work

[Jamie / Suresh / Kiwamu] 

The in-vac work is ongoing.

Before we run out our energy we are posting this entry to briefly report the current status.

- (done) BS earthquake stop adjustment.

- (done) PRM earthquake stop adjustment

- (done) MC spot position check => They are almost the same within 10 %.

- (done) Injection and alignment of the ABSL laser to make the beam brighter in the vertex region.

- (done) POY => We repositioned an in-vac steering mirror to get the POY beam hitting the center of the steering mirror.

                           It's now coming out from the chamber.

(done)  IPANG => realigned two mirrors on the ETMY table to get the IPANG out from the chamber. Now it's reaching the ETMY optical table.

                             It needs a final touch before we pump down.  We revisited it later in the night after realigning the IFO and it is well aligned now.

- (done)  POP => We have aligned the ABSL laser injected from the AS port to reach the REFL camera.  We turned it up to max power of 300mW and used it as a substitute for the PRC beam.

                         Even this was not enough to see anything in the POP beam path after the PR2 (tip-tilt).  So we used a green beam from the Y-arm as a guide of the POP beam path because the ABSL (POP) beam was too dim to work with.

                         We placed a lens and a CCD camera to detect the green and then blocked the Y-green.  It was then possible to see the ABSL-POP beam in the CCD camera.  The lens and the CCD are markers for this beam. 

                         Do not remove these markers unless absolutely necessary.

-(done) POX => We located the ABSL (mimicking POX) beam on the POXM1 mirror and adjusted the mirror to ensure that the beam exits at the right height and a convenient location on the POX table. 

- (0%) OSEM mid-range adjustment

- (0%) IPPOS

- (0%) oplev re-alignment

  5251   Wed Aug 17 02:48:56 2011 kiwamuUpdateRF SystemRe: AM in the PM

[Keiko / Suresh / Anamaria / Kiwamu]

 The AM components do exist also on the beam after the EOM.

The peaks were found at 11, 29 and 55 MHz, where the PM are supposed to be imposed.

Suresh and Keiko minimized them by rotating the HWP, which is in front of the EOM.

Also Anamaria and I tried minimizing them by adjusting the EOM crystal alignment.

However everytime after we minimized the AM peaks, they grew back in a time scale of ~ 1 min.

Potentially it could be a problem of the HWP and/or EOM alignment.

Since we wanted to proceed the in-vac work anyways, we stopped investigating it and decided to postpone it for tomorrow.

We again adjusted the incident power to 20 mW.

 

-- P.S.

 The incident power going to MC went down to 7 mW for some reasons. This was found after ~ 6 hours from our works on the PSL table.

We haven't touched anything on the PSL table since the daytime work.

Possibly the angle of the HWP is drifting (why?) and changed the amount of the P-polarizing beam power.

Suresh locked the angles of two HWPs, which are the one just after the EOM and the one after the attenuation PBS.

Quote from #5249

So first check is beam out of PSL EOM, to make sure the input beam is aligned to the crystal axis and is not giving AM modulation in adition to PM.

 

  5250   Tue Aug 16 17:09:55 2011 JenneUpdateGeneraltoday's work to do

Quote:

  + Rotate the SRM tower to get the SRMI fringes on the AS CCD camera.

        => This is because the required amount of the YAW correction on SRM is currently beyond the range of the DC bias.

 Kiwamu aligned things for me, and I rotated the SRM tower so that the reflected beam was pretty much totally overlapping the incident beam.  The SRC was aligned to make sure things were good.  Now the DC bias for SRM Yaw is ~1.4, so we're totally good. 

To rotate SRM, Jamie had the idea of using 2 screws so I could push the tower on one side, and back off the screw an equal amount on the other side and push the tower to be touching both screws again, to ensure that I was rotating about the center of the tower and wasn't introducing any Pos action. 

While I was at it, I also moved the OSEM connector tower back to its normal place on the table, so it's not in the way of oplev beams.  It had been moved previously to accommodate ITMY near the door.

  5249   Tue Aug 16 16:59:20 2011 AnamariaUpdateRF SystemAM in the PM

Kiwamu, Keiko, Anamaria

Looking at the I and Q signals coming from REFL11 and REFL55 we saw large offsets, which would mean we have amplitude modulation, especially at 11MHz. We checked the PD themselves with RF spectrum analyzer, and at their frequencies we see stationary peaks (even if we look only at direct reflection from PRM). We changed the attenuation of the PSL EOM, and saw the peak go down. So first check is beam out of PSL EOM, to make sure the input beam is aligned to the crystal axis and is not giving AM modulation in adition to PM.

  5248   Tue Aug 16 11:49:17 2011 jamie, jenneUpdateGeneraltoday's work to do

>If necessary steer ETMs and ITMs to make the X and Y green beam flashing.

Green is now flashing in both X and Y arms

>Open the IOO and OMC chamber and lock MC.

Open, and cover in place. MC is flashing and locking.

 

  5247   Tue Aug 16 10:59:06 2011 jamieUpdateSUSSUS update

Data taken from: 997530498+120

Things are actually looking ok at the moment.  "Badness" (cond(B)) is below 6 for all optics.

  • We don't have results from PRM since its spectra looked bad, as if it's being clamped by the earthquake stops.
  • The SRM matrix definitely looks the nicest, followed by ITMX.  All the other matrices have some abnormally high or low elements.
  • cond(B) for ETMY is better than that for SRM, even though the ETMY matrix doesn't look as nice.  Does this mean that cond(B) is not necessarily the best figure of merit, or is there something else that our naive expectation for the matrix doesn't catch?

We still need to go through and adjust all the OSEM ranges once the IFO is aligned and we know what our DC biases are.  We'll repeat this one last time after that.

TM   M cond(B)
BS  BS.png       pit     yaw     pos     side    butt
UL    1.456   0.770   0.296   0.303   1.035 
UR    0.285  -1.230   1.773  -0.077  -0.945 
LR   -1.715  -0.340   1.704  -0.115   0.951 
LL   -0.544   1.660   0.227   0.265  -1.070 
SD    0.612   0.275  -3.459   1.000   0.046
5.61948
SRM  SRM.png       pit     yaw     pos     side    butt
UL    0.891   1.125   0.950  -0.077   0.984 
UR    0.934  -0.875   0.987  -0.011  -0.933 
LR   -1.066  -1.020   1.050   0.010   1.084 
LL   -1.109   0.980   1.013  -0.056  -0.999 
SD    0.257  -0.021   0.304   1.000   0.006 
4.0291
ITMX  ITMX.png       pit     yaw     pos     side    butt
UL    0.436   1.035   1.042  -0.068   0.728 
UR    0.855  -0.965   1.137  -0.211  -0.969 
LR   -1.145  -1.228   0.958  -0.263   1.224 
LL   -1.564   0.772   0.863  -0.120  -1.079 
SD   -0.522  -0.763   2.495   1.000  -0.156
4.55925
ITMY  ITMY.png       pit     yaw     pos     side    butt
UL    1.375   0.095   1.245  -0.058   0.989 
UR   -0.411   1.778   0.975  -0.022  -1.065 
LR   -2.000  -0.222   0.755   0.006   1.001 
LL   -0.214  -1.905   1.025  -0.030  -0.945 
SD    0.011  -0.686   0.804   1.000   0.240 
4.14139
ETMX  ETMX.png       pit     yaw     pos     side    butt
UL    0.714   0.191   1.640   0.404   1.052 
UR    0.197  -1.809   1.758  -0.120  -1.133 
LR   -1.803  -1.889   0.360  -0.109   0.913 
LL   -1.286   0.111   0.242   0.415  -0.902 
SD    1.823  -3.738  -0.714   1.000  -0.130 
5.19482
ETMY  ETMY.png       pit     yaw     pos     side    butt
UL    1.104   0.384   1.417   0.351   1.013 
UR   -0.287  -1.501   1.310  -0.074  -1.032 
LR   -2.000   0.115   0.583  -0.045   0.777 
LL   -0.609   2.000   0.690   0.380  -1.179 
SD    0.043  -0.742  -0.941   1.000   0.338 
3.57032

 

  5246   Tue Aug 16 04:50:17 2011 kiwamuUpdateSUSfree swinging again

 Since Suresh and I changed the DC biases on most of the suspension, the free swingning spectra will be different from the past.

- -

EXcited ETMX ETMY ITMX ITMY PRM SRM BS

Tue Aug 16 04:48:02 PDT 2011
997530498

  5245   Tue Aug 16 04:28:09 2011 kiwamuUpdateGeneraltoday's work to do

The main goal of today is to extract the pick-off beams

Today's menu :

  + If necessary steer ETMs and ITMs to make the X and Y green beam flashing.

  + Open the IOO and OMC chamber and lock MC.

       => cover the place of the access connector by a large piece of aluminum foil. It will give a robust lock of MC.

  + Check the beam pointing down to Y arm by looking at the ETMY face camera.

        => If it's necessary align PZT1 and PZT2 from EPICS to make the IR beam flashing in the X arm.

  + Align BS and let the beam hit the center of ETMX to make the X arm flashing.

        => These alignment procedure will automatically gives us the MICH fringes on the AS CCD camera.

  + Rotate the SRM tower to get the SRMI fringes on the AS CCD camera.

        => This is because the required amount of the YAW correction on SRM is currently beyond the range of the DC bias.

  + Align PRM to get the PRMI fringes on AS CCD camera. Also make sure the beam comes back to the REFL CCD camera.

  + Lock the PRC to make POP/POX/POY bright enough.

     => Probably the REFL11 RFPD needs more power. To increase the power, just rotate the HWP, which is sitting before the RFPDs on the AS table.

     => If the signal on AS55 is too small, we can use REFL11_Q or REFL55 for the MICH lock.

  + OR inject and align the ABSL laser to make those pick-off beams bright enough.

     => This case we don't have to make the DRMI exactly on the resonance, what we need is just the DRMI flashing.

  + Align necessary optics for those pick-off beams.

     => In our definition (aLIGO definition) POP is the beam propagating from BS to PRM. Don't be confused by another one, which propagates from PRM to BS.

  + Steer two mirrors on the ETMY table for aligning IPANG. Also, steer some mirrors on the BS table for IPPOS.

    => IPANG has already reached the ETMY table, so ideally we don't have to steer a mirror on the BS table.

    => IPPOS/ANG are very visible with a sensor card.

  + Align some oplevs.

     => This work is relatively independent from the other tasks. Steve can take care of it.

  + Adjust the OSEM to their mid-range.

     => This work can be done anytime, but everytime we have to recover the alignment of the interferometer.

  5244   Tue Aug 16 04:25:34 2011 Suresh, KiwamuUpdateSUSalignment of MC output to Y-arm using PZTs

We did several things today+night.  The final goal was to lock the PRC so that we could obtain the POX, POY and POP beams.  However there were large number of steps to get there.

1) We moved the ITMY into its place and balanced the table

2) We then aligned the Y-arm cavity to the green beam which was set up as a reference before we moved the ETMY and ITMY to adjust the OSEMS.  We had the green flashing in Y-arm

3) We checked the beam position on PR2. It was okay. This confirmed that we were ready to send the beam onto the Y arm.

4) We then roughly aligned the IR beam on ETMY where Jamie had placed an Al foil with a hole.  We got the arm flashing in both IR and green. 

5) We used the PZTs to make the green and IR beams co-incident and flashing in the Y arm.  This completed the alignment of the IR beam into the Y-arm.

6) The IPPO (pick-off) window had to be repositioned to avoid clipping.  The IPANG beam was aligned such that it exits the ETMY chamber onto the ETMY table.  It can now be easily sent to the IPANG QPD.

7) Then BS was aligned to direct the IR beam into the X-arm and had the X-arm flashing.  It had already been aligned to its green.

8) It was now the turn of the SRC.  The beam spots on all the SRC related optics were off centered.  We aligned all the optics in the AS path to get the AS beam on to the AP table.

9) The AS beam was very faint so we repositioned the AS camera to the place intended for AS11 PD, since there was a brighter beam available there. 

10) We could then obtain reflections from ITMY, ITMX and PRM at the AS camera. 

11) Problems:

      a) ITMY osems need to be readjusted to make sure that they are in mid-range.  Several are out of range and so the damping is not effective.

      b) When we tried to align SRC the yaw OSEM had to be pushed to its full range.  We therefore have to turn the SRM tower to get it back into range.

 12)  We stopped here since moving the SRM is not something to be attempted at the end of a rather long day. Kiwamu is posting a plan for the rest of the day.

  5243   Mon Aug 15 21:43:29 2011 Anamaria and KeikoSummaryLockingcentral part ifo locking project

 REFL33 and REFL165 cables were connected from the AP table to the rack.  Cables on the rack for REFL33I, 33Q, 165I, 165Q ports were connected, too. Connections were confirmed by the data viewer. Two SMA cables which will be used for the two PDs on the AP tabl were built. We will be able to place the two PDs tomorrow. The beamsplitters to split the laser to REFL33 and REFL165 ports were mounted and ready to be placed.

  5242   Mon Aug 15 17:38:07 2011 jamieUpdateGeneralFoil aperture placed in front of ETMY

We have placed a foil aperture in front of ETMY, to aid in aligning the Y-arm, and then the PRC.  It obviously needs to be removed before we close up.

  5241   Mon Aug 15 17:36:10 2011 jamieUpdateSUSStrangeness with ETMY (was: Monday SUS update)

For some reason ETMY has changed a lot.  Not only does it now have the worst "badness" (B matrix condition number) at ~10, but the frequency of all the modes have shifted, some considerably.  I did accidentally bump the optic when Jenne and I were adjusting the OSEMs last week, but I didn't think it was that much.  The only thing I can think of that would cause the modes to move so much is that the optic has been somehow reseated in it's suspension.  I really don't know how that would have happened, though.

Jenne and I went in to investigate ETMY, to see if we could see anything obviously wrong.  Everything looks to be ok.  The magnets are all well centered in the OSEMs, and the PDMon levels look ok.

We rechecked the balance of the table, and tweaked it a bit to make it more level.  We then tweaked the OSEMs again to put them back in the center of their range.  We also checked the response by using the lockin method to check the response to POS and SIDE drive in each of the OSEMs (we want large POS response and minimal SIDE response).  Everything looked ok.

We're going to take another freeswing measurement and see how things look now.  If there are any suggestions what should be done (if anything), about the shifted modes, please let us know.

  5240   Mon Aug 15 17:23:55 2011 jamieUpdateSUSfreeswing script updated

I have updated the freeswing scripts, combining all of them into a single script that takes arguments to specify the optic to kick:

pianosa:SUS 0> ./freeswing
usage: freeswing SET
usage: freeswing OPTIC [OPTIC ...]

Kick and free-swing suspended optics.
Specify optics (i.e. 'MC1', 'ITMY') or a set:
'all' = (MC1 MC2 MC3 ETMX ETMY ITMX ITMY PRM SRM BS)
'ifo' = (ETMX ETMY ITMX ITMY PRM SRM BS)
'mc'  = (MC1 MC2 MC3)
pianosa:SUS 0>

I have removed all of the old scripts, and committed the new one to the SVN.

  5239   Mon Aug 15 14:10:56 2011 JenneUpdateSUSMonday SUS update

The moral of the story here is that none of the suspensions are overwhelmingly awesome, but most of them will be fine if we leave them as-is.

SUS DoF Plot Input Matrix "BADness" (1==good)
ITMX inMatDiag.png       pit     yaw     pos     side    butt
UL    0.438   1.019   1.050  -0.059   0.717 
UR    0.828  -0.981   1.128  -0.215  -0.956 
LR   -1.172  -1.201   0.950  -0.275   1.241 
LL   -1.562   0.799   0.872  -0.120  -1.087 
SD   -0.579  -0.847   2.539   1.000  -0.170 

 
4.68597
 
 ITMY  inMatDiag.png        pit     yaw     pos     side    butt
UL    1.157   0.185   1.188  -0.109   0.922 
UR    0.020  -1.815   0.745  -0.051  -0.970 
LR   -1.980  -0.090   0.812  -0.024   1.158 
LL   -0.843   1.910   1.255  -0.082  -0.949 
SD   -0.958   1.080   1.859   1.000   0.325  
4.82756
ETMX  inMatDiag.png        pit     yaw     pos     side    butt
UL    0.338   0.476   1.609   0.316   1.046  
UR    0.274  -1.524   1.796  -0.069  -1.180  
LR   -1.726  -1.565   0.391  -0.100   0.938  
LL   -1.662   0.435   0.204   0.286  -0.836  
SD    0.996  -2.629  -0.999   1.000  -0.111
 
 
 4.32072
 ETMY  inMatDiag.png        pit     yaw     pos     side    butt
UL    1.123   0.456   1.812   0.231   0.936 
UR   -0.198  -1.489   0.492  -0.096  -1.098 
LR   -2.000   0.055   0.188  -0.052   0.764 
LL   -0.679   2.000   1.508   0.275  -1.201 
SD    0.180  -0.591   3.355   1.000   0.200  
 10.643
 BS  inMatDiag.png        pit     yaw     pos     side    butt
UL    1.575   0.697   0.230   0.294   1.045 
UR    0.163  -1.303   1.829  -0.133  -0.958 
LR   -1.837  -0.308   1.770  -0.171   0.944 
LL   -0.425   1.692   0.171   0.257  -1.053 
SD    0.769   0.345  -3.380   1.000   0.058 
6.111
 
 PRM  inMatDiag.png        pit     yaw     pos     side    butt
UL    0.597   1.553   2.000  -0.469   1.229  
UR    1.304  -0.447   0.383  -0.043  -0.734  
LR   -0.696  -1.048  -0.277   0.109   0.687  
LL   -1.403   0.952   1.340  -0.317  -1.350  
SD    0.518  -1.125  -1.161   1.000   0.394  

 
 8.43363
SRM inMatDiag.png       pit     yaw     pos     side    butt
UL    0.831   1.039   1.153  -0.140   1.065 
UR    1.071  -0.961   1.104  -0.057  -1.061 
LR   -0.929  -0.946   0.847  -0.035   0.837 
LL   -1.169   1.054   0.896  -0.118  -1.037 
SD    0.193  -0.033   1.797   1.000   0.045 
 4.17396

 

  5238   Mon Aug 15 14:07:39 2011 kiwamuUpdateIOORe: MC misaligned a lot

The leveling was still okay. The MC mirrors were realigned and now they all are fine.

We will go ahead for the vertex alignment and extraction of the pick-off beams.

 

Here is a summary of the spot measurement.

    Feb 26 2011      May 08 2011 Aug 2 2011  Aug 10 2011 (in air) [NEW!!] Aug 14 2011 (in air)
MC1 pit [mm]   1.6   1.9  1.93 -0.858 -0.2
MC2 pit [mm]   6.4   9.0 9.03 -0.844 -0.8
MC3 pit [mm]   1.4   2.0 2.01 -1.03 -0.1
MC1 yaw [mm]   -1.5   -1.7 -1.72 -0.847 -1.1
MC2 yaw [mm]   1.0   0.2 0.178 0.582 0.6
MC3 yaw [mm]   -1.3   -1.9 -1.87 -1.06 -1.1

 

Quote from #5236

Anyways we should check the leveling of the IOO table and the spot positions on the MC mirrors again to make sure.

 

  5237   Mon Aug 15 13:16:50 2011 JenneUpdateSUSRe: ETMY hopefully good again

Quote:

I guess the ETMY suspension is still fine. Their OSEM DC voltage and the free swinging spectra look healthy.

It could be a failure in the initial guess for fitting.

Quote from #5216

I'm no longer convinced that ETMY is healthy. I can't fit the peaks to get the input matrix.

 Turns out I was missing a critical step in the process...running makeSUSspectra.m  After I do that, everything is back under control, and ETMY looks fine. 

I'm almost done doing the peak-fitting and matrix inversion for all optics.

  5236   Mon Aug 15 10:58:52 2011 kiwamuUpdateIOOMC misaligned a lot

This morning Steve and I opened the doors on the IOO and OMC chamber to let the IR beam go to MC.

And found the MC flashing is way far from TEM00, there were very higher order modes.

The MC suspensions were realigned based on an assumption that the incident beam didn't change recently.

Anyways we should check the leveling of the IOO table and the spot positions on the MC mirrors again to make sure.

  5235   Mon Aug 15 10:02:27 2011 kiwamuUpdateSUSBS OSEM adjustment done

[Suresh / Kiwamu]

 Adjustment of the OSEMs on BS has been done.

All the bad suspensions (#5176) has been adjusted. They are waiting for the matrix inversion test.

  5234   Sun Aug 14 22:48:37 2011 kiwamuUpdateSUSfree swinging again

Excited all optics
Sun Aug 14 20:22:33 PDT 2011
997413768

  5233   Sun Aug 14 20:04:40 2011 Keiko, Anamaria, Jenne, and KiwamuSummaryLockingcentral part ifo locking plan
GOAL : To lock the central part of ifo

Here is the plan:

Mon - assemble all the cables from PDs and mixers, and check the CDS channels. Prepare the beamsplitters.

Tue - The current paths to REFL11 and REFL55 will be modified to the four paths to REFL11, 33, 55, 165. And the PDs will be placed.
Wed, Thu - during waiting for the ifo available with vacuum, help aligning the POP, POX, POY. In parallel, a simulation to find the PRC length SRC 
length tolerance will be proceeded.

Fri - When the ifo becomes available with vacuum, the sensing signals by 3-f scheme will be obtained with proper demodulation phases.

Sat - Try to lock the central part of the ifo with the new 3-f signals.
  5232   Sun Aug 14 19:06:50 2011 JenneUpdateelogelog dead. Brought back to life

like the subject says...

  5231   Sun Aug 14 17:47:39 2011 NicoleSummarySUSTT Shaking Today and Hopefully More?

Quote:

As reported in my  previous entry of TT supsension bode plots, I found that my experimental data had what appears to be very noise peaks above 20 Hz (as mentioned earlier, the peak at 22 Hz is likely due to vertical coupling, as 22 Hz is the resonant frequency of the cantilever blades). This is very unusual and needs to be explored further. I would like to vertically-shake the TTs to obtain more data on possible coupling. However, I am leaving on Monday and will not return until Thursday (day of SURF talks). I am leaving campus Friday afternoon or so. I would may need some help coming up with an assembly plan/assembling set-up for vertical shaking (if it is possible to do so in such a limited time frame).

 

Today I wanted to see if the "noisy peaks" above 30 Hz were due to EM noise coupling. I tested this hypothesis today, seeing if EM fields generated by the coil at higher frequencies were injecting noise into my transfer function measurements. I found that the "noisy peaks" above 30 Hz are NOT DUE TO EM NOISE COUPLING. I am very curious as to what is causing the high peaks (possibly coupling from other degrees of freedom)?

 

 emnoise_and_flex.jpg

 I have been redoing the noise test multiple times today. Here is the best plot that I got

noisetest.jpg

  5230   Sun Aug 14 15:37:39 2011 JennyUpdatePSLTemperature steps and slow actuator railing

Quote:

Below are some plots from dataviewer of temperature-step data taken over the past 32 hours. (They show minute trends). I am looking at the thermal coupling from the can surrounding the reference cavity on the PSL table to the cavity itself, and trying to measure the cavity temperature response via the control signal sent to heat the NPRO laser, which is locked to the cavity.

Picture_6.png

Picture_7.png

  • Top left: out-of-loop temperature sensor on can surrounding ref cav (RCTEMP)
  • Top right: control signal sent to slow drive of laser (laser heater), which is supposed to follow the cavity temperature (TMP_OUTPUT)
  • Bottom left: in-loop can temperature sensors (MINCOMEAS)
  • Bottom right: room temperature reading (RMTEMP)

 

I stepped the temperature set point from 35 to 36 deg. C for the can at 12:30am last night. Then I waited to see the cavity temperature change and the slow actuator (laser heater: TMP_OUTPUT) follow that change.

I was a bit worried about the oscillations that were occuring in the TMP_OUTPUT signal even long after this temperature step was made, but I figured that they were simply room-temperature changes propagating into the cavity, since they seemed to have a similar pattern to the room-temperature variations, and since it is clear that the out-of-loop temperature sensor on the can (RCTEMP) experiences variations, even when the in-loop sensors are recording no variation.

At 8:46pm tonight I stepped the temperature down 2 degrees to 34 deg. C. The step had a clear effect on TMP_OUTPUT. The voltage to the heater dropped and eventually railed at its lowest output. I'm worried that the loop is unstable, although I haven't ruled out other possibilities, such as that a 2 deg. C temperature step is too large for the loop. I will investigate further in the morning.

 The lock was lost when I came in around noon today to check on it. The slow actuator was still railing.

1) I got lock back for a few minutes, by varying the laser temperature set point manually. TMP_OUTPUT was still reading -30000 cts (minimum allowed) and the transmission was not as high as it had been.

2) I toggled the second filter button off. The TMP_OUTPUT started rising up to ~2000 cts. I then toggled the second filter back on, and TMP_OUTPUT jumped the positive maximum number of counts allowed.

3) I lost the lock again. I turned off the digital output to the slow actuator.

4) I have so far failed at getting the lock back. My main problem is that when the BNC cable to the slow port is plugged in, even when I'm not sending anything to that port, it makes it so that changing the temperature set point manually has almost no effect on the transmission (it looks as though changing the setpoint is not actually changing the temperature, because the monitor shows the same higher order mode even when with +-degree temperature setpoint changes).

  5229   Sun Aug 14 13:57:52 2011 NicoleSummarySUSTT Optimization Curves

Quote:

Quote:

1) Drawing has the dimensions => You can calculate the volume => You can calculate the mass
Complicated structure can be ignored. We need a rough estimation.

2) Your restoring force can have two terms:
- one comes from the spring constant k
- the other from the gravity

 

 Thank you.

 

The wire used to suspend the EDCs is tungsten?

To verify, for my model, the EDC will be the mass of all four dampers or a single damper? The length of the wire used to suspend the EDC will be the combined length of 4 wires or length of  a single wire?

 

Taking into account the densities for each material (specific material of each component was listed, so I looked up the densities), and trying my best to approximate the volumes of each component, I have determined

the mass of the mirror + mirror holder to be ~100 g and the mass of a single EDC to be ~19 g

 I am thinking that perhaps my mass estimations were off? The model that I have used fits the data better than the model that I have made (changing the masses to fit my estimations of the values)

FLEX_exp_mod_comp.jpg

  5228   Sun Aug 14 04:12:37 2011 JennyUpdatePSLTemperature steps and slow actuator railing

Below are some plots from dataviewer of temperature-step data taken over the past 32 hours. (They show minute trends). I am looking at the thermal coupling from the can surrounding the reference cavity on the PSL table to the cavity itself, and trying to measure the cavity temperature response via the control signal sent to heat the NPRO laser, which is locked to the cavity.

Picture_6.png

Picture_7.png

  • Top left: out-of-loop temperature sensor on can surrounding ref cav (RCTEMP)
  • Top right: control signal sent to slow drive of laser (laser heater), which is supposed to follow the cavity temperature (TMP_OUTPUT)
  • Bottom left: in-loop can temperature sensors (MINCOMEAS)
  • Bottom right: room temperature reading (RMTEMP)

 

I stepped the temperature set point from 35 to 36 deg. C for the can at 12:30am last night. Then I waited to see the cavity temperature change and the slow actuator (laser heater: TMP_OUTPUT) follow that change.

I was a bit worried about the oscillations that were occuring in the TMP_OUTPUT signal even long after this temperature step was made, but I figured that they were simply room-temperature changes propagating into the cavity, since they seemed to have a similar pattern to the room-temperature variations, and since it is clear that the out-of-loop temperature sensor on the can (RCTEMP) experiences variations, even when the in-loop sensors are recording no variation.

At 8:46pm tonight I stepped the temperature down 2 degrees to 34 deg. C. The step had a clear effect on TMP_OUTPUT. The voltage to the heater dropped and eventually railed at its lowest output. I'm worried that the loop is unstable, although I haven't ruled out other possibilities, such as that a 2 deg. C temperature step is too large for the loop. I will investigate further in the morning.

  5227   Sun Aug 14 00:26:51 2011 NicoleSummarySUSTT Optimization Curves

Quote:

1) Drawing has the dimensions => You can calculate the volume => You can calculate the mass
Complicated structure can be ignored. We need a rough estimation.

2) Your restoring force can have two terms:
- one comes from the spring constant k
- the other from the gravity

 

 Thank you.

 

The wire used to suspend the EDCs is tungsten?

To verify, for my model, the EDC will be the mass of all four dampers or a single damper? The length of the wire used to suspend the EDC will be the combined length of 4 wires or length of  a single wire?

 

Taking into account the densities for each material (specific material of each component was listed, so I looked up the densities), and trying my best to approximate the volumes of each component, I have determined

the mass of the mirror + mirror holder to be ~100 g and the mass of a single EDC to be ~19 g

  5226   Sat Aug 13 21:48:17 2011 KojiSummarySUSTT Optimization Curves

1) Drawing has the dimensions => You can calculate the volume => You can calculate the mass
Complicated structure can be ignored. We need a rough estimation.

2) Your restoring force can have two terms:
- one comes from the spring constant k
- the other from the gravity

 

  5225   Sat Aug 13 21:15:47 2011 NicoleSummarySUSTT Optimization Curves

Quote:

What are the parameters you are using? As you have the drawings of the components, you can calculate the masses of the objects.

 

Reducing the ECD resonance from 10Hz->6Hz looks nice.

The resonant freq of the ECDs are not (fully) determined by the gravitational energy but have the contribution of the elastic energy of the wire.

Q1: How much is the res freq of the ECDs if the freq is completely determined by the grav energy? (i.e. the case of using much thinner wires)

Q2: How thin should the wires be?

 The drawings do not have the masses of the objects.

 

For the resonant frequency:

Instead of sqrt (g/l) would the numerator in the square root be[ g + (energy stored in wire)/(mass of damper)] ?

 

  5224   Sat Aug 13 19:08:01 2011 KojiSummarySUSTT Optimization Curves

What are the parameters you are using? As you have the drawings of the components, you can calculate the masses of the objects.

 

Reducing the ECD resonance from 10Hz->6Hz looks nice.

The resonant freq of the ECDs are not (fully) determined by the gravitational energy but have the contribution of the elastic energy of the wire.

Q1: How much is the res freq of the ECDs if the freq is completely determined by the grav energy? (i.e. the case of using much thinner wires)

Q2: How thin should the wires be?

  5223   Sat Aug 13 15:47:47 2011 NicoleSummarySUSTT Optimization Curves

Using my Matlab model of the flexibly-supported eddy current damping system, I have changed parameters to see if/how the TTs can be optimized in isolation. As I found earlier, posted in my bode plot entry, there is only a limited region where the flexibly-supported system provides better isolation than the rigidly-supported system.

 

Here is what I have found, where \gamma is the scale factor of the magnetic strength (proportional to magnetic strength), \beta is the scale factor of the current damper mass (estimated by attempting to fit my model to the experimental data), and \alpha is the scale factor of the current resonant frequency of the dampers.

magstrength1.jpgdampermass.jpg resfreq.jpg

 

Here are my commentaries on these plots. If you have any commentaries, it would be very helpful, as I would like to incorporate this information in my powerpoint presentation.

It seems as if the TT suspensions are already optimized?

It may be difficult to lower the resonant frequency of the dampers because that would mean changing the lengths of the EDC suspensions). Also, it appears that a rather drastic reduction (at most 0.6*current EDC resonant frequency --> reduction from about 10 Hz to 6 Hz or less) is required . Using the calculation that the resonant frequency is sqrt(g/length), for my single-suspended EDC model, this means increasing the wire length to nearly 3 x its current value. I'm not sure how this would translate to four EDCs...

The amplification at resonance caused by increasing the magnet strength almost offsets the isolation benefits of increasing magnet strength. From my modeling, it appears that the magnet strength may be very close (if not already at) isolation optimization.

 

Lowering the mass to 0.2 the current mass may be impractical. It seems as if the benefits of lowering the mass only occur when the mass is reduced by a factor of 0.2 (maybe 0.4)

 

  5222   Sat Aug 13 15:40:38 2011 NicoleSummarySUSTT Shaking Today and Hopefully More?

As reported in my  previous entry of TT supsension bode plots, I found that my experimental data had what appears to be very noise peaks above 20 Hz (as mentioned earlier, the peak at 22 Hz is likely due to vertical coupling, as 22 Hz is the resonant frequency of the cantilever blades). This is very unusual and needs to be explored further. I would like to vertically-shake the TTs to obtain more data on possible coupling. However, I am leaving on Monday and will not return until Thursday (day of SURF talks). I am leaving campus Friday afternoon or so. I would may need some help coming up with an assembly plan/assembling set-up for vertical shaking (if it is possible to do so in such a limited time frame).

 

Today I wanted to see if the "noisy peaks" above 30 Hz were due to EM noise coupling. I tested this hypothesis today, seeing if EM fields generated by the coil at higher frequencies were injecting noise into my transfer function measurements. I found that the "noisy peaks" above 30 Hz are NOT DUE TO EM NOISE COUPLING. I am very curious as to what is causing the high peaks (possibly coupling from other degrees of freedom)?

 

 emnoise_and_flex.jpg

  5221   Sat Aug 13 02:31:42 2011 kiwamuUpdateSUSRe: ETMY hopefully good again

I guess the ETMY suspension is still fine. Their OSEM DC voltage and the free swinging spectra look healthy.

It could be a failure in the initial guess for fitting.

Quote from #5216

I'm no longer convinced that ETMY is healthy. I can't fit the peaks to get the input matrix.

  5220   Sat Aug 13 02:11:33 2011 kiwamuUpdateSUSfree swinging again

I am leaving all of the suspensions free swinging. They will automatically recover after 5 hours from now.

--
Excited all optics
Sat Aug 13 02:08:07 PDT 2011
997261703
--

FYI : I ran a combination of two scripts:   ./freeswing && ./opticshutdown

  5219   Sat Aug 13 01:54:18 2011 kiwamuUpdateSUSPRM OSEM adjustment part II

Adjustment of the PRM OSEMs are done. The coils turned out to be healthy.

The malfunction was fixed. It was because the UL OSEM was too deeply inserted and barely touching the AR surface of the mirror.

 

(OSEM adjustment)

 + Excited POS at 6.5 Hz with an amplitude of 3000 cnts by the LOCKIN oscillator.

 + Looked at the signal of each sensor in frequency domain.

 + Maximized the excitation peak for each of the four face OSEMs by rotating them.

 + Minimized the excitation peak in the SIDE signal by rotating it.

 + Adjusted the OSEM translational position so that they are in the midpoint of the OSEM range.

 

(POS sensitivity check)

From the view point of the matrix inversion, one thing we want to have is the equally sensitive face sensors and insensitive SIDE OSEM to the POS motion.

To check the success level of today's PRM adjusment, I ran swept sine measurements to take the transfer function from POS to each sensor.

The plots below are the results.  The first figure is the one measured before the adjustment and the second plot is the one after the adjustment.

As shown in the plot, before the adjustment the sensitivity of OSEMs were very different and the SIDE OSEM is quite sensitive to the POS motion.

So PRM used be in an extremely bad situation.

After the adjustment, the plot became much better.

The four face sensors have almost the same sensitivity (within factor of 3) and the SIDE is quite insensitive to the POS motion.

before.png after.png

Quote from #5203

 We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.

The UL and LL coils are not working correctly, the forces are weak.

  5218   Sat Aug 13 01:52:07 2011 YoichiUpdateLSCFeed forward delay
Yoichi, Koji

While I was testing the feed forward cancellation, I noticed that the
cancellation was not perfect.
The test I did was the following.
I injected the same signal to both DARM and MICH feedback filters.
This was done by injecting a signal into the excitation point of
the ASDC PD, then changing the input matrix elements so that the signal
goes to both DARM and MICH.
Then in the FFC, MICH signal was fed forward to DARM by the gain of -1.
Ideally, this should completely eliminate the DARM FB signal.
In reality, it did not.

The first PDF compares the spectrum of the injected noise (white noise,
red curve) with the spectrum of the signal after the FFC (blue curve).
At higher frequencies, the cancellation becomes poor.
It suggests that this is caused by some delay in the FFC.
I also took a transfer function from the injection point to the signal
after the FFC (second attachment).
I fitted the measured TF with a theoretical formula of
1-exp(-i*dt*f),
where dt is the time delay and f is the frequency.
The fitting is very good, and I got dt = 0.8msec ~ 13 samples for 16kHz.
13 samples is something very large.

The cause of the delay was suspected to be the shared memory communication
between different processes.
I moved all the FFC blocks to c1lsc.mdl.
Then the cancellation becomes perfect. The signal after the FFC is
completely zero, so I couldn't even make a TF measurement.

This results suggest that a large delay of 13 samples is induced
when you use shared memory to send signals round trip.
We should make simpler models, just passing signals back and forth
via shared memory, dolphin network or GE FANAC RFM to check the
delays more precisely.

For the moment, the FCC is included in the c1lsc model.
The MEDM screens were modified to account for this change.
c1ffc is stopped and removed from rtsystab.
Attachment 1: Spe1.pdf
Spe1.pdf
Attachment 2: TFFitting.png
TFFitting.png
  5217   Fri Aug 12 20:33:57 2011 DmassSummaryPSLNPRO PDH-Locked to Ref Cav

To aid Jenny's valiant attempt to finish her SURF project, I did some things with the front end system over the last couple days, largely tricking Jamie into doing things for me lest I ruin the 40m RCG system. Several tribulations have been omitted.

We stole a channel in the frontend, in the proccess:

  1. Modified the C1GFD simulink model (now analog) to be "ADC -> TMP -> DAC" where TMP is a filter bank
    • C1GFD_TMP.adl (in /opt/rtcds/caltech/c1/medm/c1gfd) is the relevant part which connects the ADC to the DAC in the frontend
  2. Confirmed that the ADC was working by putting a signal in and seeing it in the frontend
  3. Could not get a signal out of the anti aliasing board
  4. Looked sad until Kiwamu found a breakout board for the SCSI cable coming from the DAC
  5. Used SR560 to buffer DAC output
    • drove a triangle wave with AWG into the TMP EXC channel (100 counts 1 Hz) and looked at it after the ~25 ft of BNC cable running between the DAC and the NRPO driver
    • wave looked funny (not like a triangle wave), maybe the DAC is not meant to push a signal so far, so added buffer
  6. Took the control signal going to the fast input of the NPRO driver (using the 500 Ohm SR560 output - see Jenny's diagram) and put it into the anti aliasing board of the ADC
  7. Added switchable integrator to filter bank with Foton
    • I couldn't get the names to display in the filter bank, so I looked sad again
    • Jamie and Koji both poked at the "no name displayed" problem but had no conclusions, so I decided to ignore it
    • I confirm that when the two filters were toggled "on" that the transfer function was as expected: simple integrator with a unity gain at ~10mHz - agrees with what Foton's Bode Plot tool says it should be (see attached DTT plot)
  8. I got Jamie to manually add the two epics channels from the TMP model to the appropriate .ini file so they would be recorded
    • C1:GFD-TMP_OUTPUT  (16 Hz)
    • C1:GFD-TMP_INMON    (16 Hz)
  9. RefCav heater servo seems to still be set up, so we can use existing channels:
    • C1:PSL-FSS_RCPID_SETPOINT (temp setpoint - will do +/-1C steps about 35 C)
    • C1:PSL-FSS_MINCOMEAS (In loop temp sensor - in C)
    • C1:PSL-FSS_RCTEMP (out of loop temp sensor - in C)
    • C1:PSL-FSS_TIDALSET (Voltage to heater - rails @ +/- 2V)
  10.  Closed loop on the control signal for the NPRO driver with an integrator, saw error signal go to zero
    • Turned up gain a little bit, saw some oscillations, then turned gain down to stop them, final gain = 2
  11. Left system on for Jenny to come in and do step responses
Attachment 1: TMP_INT_TF.pdf
TMP_INT_TF.pdf
  5216   Fri Aug 12 20:28:13 2011 JenneUpdateSUSETMY hopefully good again

Quote:

[Jamie, Jenne]

We went in to have a look-see at ETMY since it looked stuck-ish.  Jamie noticed that the side magnet was pretty close to the teflon plates of the OSEM.  We rotated it a bit, and now its all better.  We also adjusted the OSEMs until their mid-ranges were happy.  The U's were a little low, and the L's were a little high, as if the optic were a bit pitched backward.  Anyhow, we checked that the table is level, and tweaked the OSEMs.  We're starting the free-swinging test now...

Excited all optics

Fri Aug 12 17:38:53 PDT 2011
997231148

 Hmmm.  I'm no longer convinced that ETMY is healthy.  I think that when I gave it a kick, it's bouncing against something.  I can't fit the peaks to get the input matrix.  I guess step 1 is to try giving it a smaller kick for the free swinging spectra.  But if the owl shift folk feel like it, they might have a look-see.

  5215   Fri Aug 12 17:37:11 2011 JenneUpdateSUSETMY hopefully good again

[Jamie, Jenne]

We went in to have a look-see at ETMY since it looked stuck-ish.  Jamie noticed that the side magnet was pretty close to the teflon plates of the OSEM.  We rotated it a bit, and now its all better.  We also adjusted the OSEMs until their mid-ranges were happy.  The U's were a little low, and the L's were a little high, as if the optic were a bit pitched backward.  Anyhow, we checked that the table is level, and tweaked the OSEMs.  We're starting the free-swinging test now...

Excited all optics

Fri Aug 12 17:38:53 PDT 2011
997231148

  5214   Fri Aug 12 17:27:49 2011 YoichiSummaryCDSToggle button for RCG
Bottom line: I made an RCG block to realize a toggle button easily.

Read on if you need such a button, or if you want to know how to
write a new RCG block with C.

-----------------
When I was making MEDM screens for FFC, I wanted to have a toggle
button to enable/disable the FFC path.
I wanted to have something like the ON/OFF buttons of the filter bank
screens, the one changes its state every time I click on it.
However, I could not find an easy way to realize that.

From MEDM, I can send a value to an EPICS channel using a "Message Button".
This value is always the same, say 1.
In the RCG model, I used a cdsEpicsMomentary block so that whenever the channel
gets 1, it stays to be 1 for a while and turns back to 0 in a second or so.
This generates a pulse of 1 when I click on a message button on a MEDM screen.
Then I needed a block to keep its internal state (0 or 1), and flips its state
whenever it receives a pulse of 1.
Since I couldn't find such a block in the current RCG library, I implemented one
using the cdsFunctionCall block. It allows you to implement a block with C code.

There is a good explanation of how to use this block in the CDS_PARTS library.
Here is basically what I did.

(1) Drag and drop thee cdsFunctionCall block to my model.

(2) In the "Block Properties", I put the following line in the Description field.
inline cdsToggle /opt/rtcds/caltech/c1/userapps/release/cds/common/src/cdsToggle.c
This means to call a function cdsToggle(), whose code is in the file indicated above.

(3) The contents of the source code is very simple.
void cdsToggle(double *in, int inSize, double *out, int outSize){
  static double x = 0;
  static double y = 0;

  if (*in != y){
    y = *in;
    if (y == 1){
      x = (x == 1) ? 0 : 1;
      *out = x;
    }
  }
}
The function prototype is always the same. *in and *out are the pointers to the arrays of doubles
for input and output signals of the block. In simuLink, the signals have to be
multiplexed so that the RCG can know how many signals are handed to or returned from the function.
In order to keep the internal state of my custom block, I used "static" keyword in the
declaration of the variables. The rest of the code should be obvious.

(4) Just compile the model as usual. The RCG will automatically include the source code and put
a call to the function in the proper place.

I made the block a library so that people can use it.
/opt/rtcds/caltech/c1/userapps/trunk/cds/common/models/cdsToggle.mdl
is the one.
For the usage of it, please have a look at
/opt/rtcds/caltech/c1/userapps/trunk/isc/c1/models/c1lsc
  5213   Fri Aug 12 17:05:22 2011 Manuel, IshwitaConfigurationPEMSTS2 Cable configuration

The WWF_M connector is the end of the STS2 seismometer orange cable and the S1 connector is the end of the gray 26-pin-cable

01050901.PDF

  5212   Fri Aug 12 16:52:49 2011 steveUpdateGeneralChamber illuminator switch

I'm looking for an ether net based  power switch for turning lights on and off for the vacuum system from MEDM screen.  This is what I found

Jamie please take a look at it.

 

 

  5211   Fri Aug 12 16:50:37 2011 YoichiConfigurationCDSFE Status screen rearranged
I rearranged the FE_STATUS.adl so that I have a space to add c1ffc in the screen.
So, please be aware that the FE monitors are no longer in their original positions
in the screen.
  5210   Fri Aug 12 16:42:51 2011 YoichiConfigurationLSCLSC Feed Forward Compensator
I've been working on adding feed forward (FF) paths to our LSC code.
So far, I've made a basic feed forward functionality connected
to the feedback path of the LSC model.

As is shown in the MEDM screen, this feed forward compensator (FFC)
takes feedback signals from several DOFs (MICH, PRCL, SRCL, CARM, XARM and YARM)
and put those signals through some filters. Then the filtered signals are
summed into the feedback signals.

There are input and output matrices to select which signal goes to which signal.

Usually, we just want to feed forward MICH to DARM. We may also want to do PRCL
to DARM and SRCL to DARM if necessary.
It is more unlikely that we use CARM for FF. But I put it there just in case.
XARM and YARM will not going to be used as is. These are place holders for future
experiments, like low frequency FF from seismic channels or something like that.
Feed forward is almost always done to DARM. But just in case we want to do some
fancier FF, like FF from PRCL to MICH, the output matrix is there to chose where
the signals will go.

I haven't really tested it because we don't have the interferometer working.
But I checked the signal flow, and it seems the model is working correctly.

=== Implementation ===
FFC is running in a separate realtime code, called c1ffc.
This is to offload c1lsc from the possible intense calculations, like adaptive stuff,
performed in the FFC in the future.
The LSC signal is passed to c1ffc via shared memory. The calculated FF signals
are passed back to c1lsc via shared memory too.
Even though FFC is in a separate realtime model, it is still conceptually a part of LSC.
So, I used top_names tag to change the names of the channels to C1:LSC-FFC_* instead of
C1:FFC-*.

In MEDM, there is an "ENABLE" button in the FF screen. Even though it is shown in the FFC
overview screen, the button itself is in the c1lsc code, so that we can disable the FFC
even when c1ffc is dead or going crazy.

=== Background ideas ===
For those of you wondering what is this feed forward thing for, I will put a brief explanation here.
Taking MICH as an example, we get the error signal for MICH from probably REFL_55Q (or AS_55Q ?).
At low frequencies, this signal properly reflects the motion of the mirrors (mostly seismic).
However, it has much worse shot noise than DARM. At higher frequencies, like above a few tens of Hz,
the error signal is dominated by the shot noise. Feeding back this signal to, say BS, means
we are shaking the BS by the shot noise, which was otherwise quiet at high frequencies.

Now, if the BS is shaken, it has some intrinsic coupling to the DARM signal.
The mechanism is that the BS motion creates some audio frequency sidebands
and this SBs reach the AS port and beat against the local oscillator to create
fake GW signals. This is called "Loop Noise Coupling".

A well known way to mitigate this problem is feed forward.
Since we know how much we are shaking the BS (because we are doing it), and
we can measure the amount of BS to DARM coupling, we can subtract out the
loop noise by feeding forward the MICH feedback signal to the DARM actuators.
In other words, the noise SBs created at the BS is canceled on the PD by the
extra SBs created at the ETMs by the feed forward.

This is what FFC is trying to do.
Attachment 1: FFCinLSC.png
FFCinLSC.png
Attachment 2: FFC.png
FFC.png
  5209   Fri Aug 12 15:46:51 2011 JenneUpdateSUSEarthquake stop procedure

According to Rana, the following is the "new" (should always have been used, but now we're going to enforce it) earthquake stop backing-off procedure:

1. Back all EQ stops away from the optic, so that it is fully free-swinging.

2. Confirm on dataviewer that the optic is truely free-swinging.

3. One at a time, slowly move the EQ stop in until it barely touches the optic.  Watch dataviewer during this procedure - as soon as the time series of the OSEMs gets a 'kink', you've just barely touched the optic.

4. Back the EQ stop off by the calculated number of turns.  No inspections, no creativity, just math.  Each EQ stop should be between 1.5m and 2.0mm away from the optic.

5. Repeat steps 3 and 4 for each EQ stop.

Note: The amount that you need to turn the screws depends on what the threads are.

FACE and TOP stops are all 1/4-20, so 1.5 turns is 1.90mm

BOTTOM stops are either #4-40 or #6-32 (depending on the suspension tower).  If #4-40, 3 turns is 1.90mm.  If #6-32, 2.5 turns is 1.98mm

  5208   Fri Aug 12 15:34:16 2011 NicoleSummarySUSBode Plot for TT Suspension

Quote:

Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)

It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?)  to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

 

 

EXPrigidvsflex.jpg

 

 

 Since the last post, I have found from the Characterization of TT data (from Jenne) that the resonant frequency of the cantilever springs for TT #4 (the model I am using) have a resonant frequency at 22 Hz. They are in fact inducing the 3rd resonance peak.

 

Here is a bode plot (CORRECTLY SCALED) comparing the rigidly-supported EDCs (model and experimental transfer functions)

RIGIDexp_model.jpg

 

Here is a bode plot comparing the flexibly-supported EDCs (model and experimental transfer functions). I have been working on this graph for FOREVER and with the set parameters, this is is close as I can get it (I've been mixing and matching parameters for well over an hour > <). I think that experimentally, the TTs have better isolation than the model because they have additional damping properties (i.e. cantilever blades that cause resonance peak at 22 Hz). Also, there may be a slight deviation because my model assumes that all four EDCs are a single EDC.

flexmodcomp.jpg

  5207   Fri Aug 12 15:16:56 2011 jamieUpdateSUStoday's SUS overview

Here's an update of the suspensions, after yesterdays in-vacuum work and OSEM tweaking:

  • PRM and ETMY are completely messed up.  The spectra are so bad that I'm not going to bother posting anything.   ETMY has extreme sensor voltages that indicate that it's maybe stuck to one of the OSEMS.  PRM voltages look nominal, so I have no idea what's going on there.
  • ITMY is much improved, but it could still use some work
  • SRM is a little worse than what it was yesterday, but we've done a lot of work on the ITMY table, such as moving ITMY suspension and rebalancing the table.
  • BS looks for some reason slightly better than it did yesterday
TM   M cond(B)
SRM SRM.png

      pit     yaw     pos     side    butt
UL    0.828   1.041   1.142  -0.135   1.057 
UR    1.061  -0.959   1.081  -0.063  -1.058 
LR   -0.939  -0.956   0.858  -0.036   0.849 
LL   -1.172   1.044   0.919  -0.108  -1.035 
SD    0.196  -0.024   1.861   1.000   0.043

4.20951
ITMY  ITMY.png       pit     yaw     pos     side    butt
UL    1.141   0.177   1.193  -0.058   0.922 
UR    0.052  -1.823   0.766  -0.031  -0.974 
LR   -1.948  -0.082   0.807  -0.013   1.147 
LL   -0.859   1.918   1.234  -0.040  -0.957 
SD   -1.916   2.178   3.558   1.000   0.635 
7.70705
BS  BS.png       pit     yaw     pos     side    butt
UL    1.589   0.694   0.182   0.302   1.042 
UR    0.157  -1.306   1.842  -0.176  -0.963 
LR   -1.843  -0.322   1.818  -0.213   0.957 
LL   -0.411   1.678   0.158   0.265  -1.038 
SD    0.754   0.298  -3.142   1.000   0.053
6.12779

 

  5206   Fri Aug 12 14:15:07 2011 NicoleSummarySUSBode Plot for TT Suspension

Here is my bode plot comparing the flexibly-supported and rigidly-supported EDCs (both with no bar)

It seems as if the rigidly-supported EDC has better isolation below 10 Hz (the mathematically-determined Matlab model predicted this...that for the same magnet strength, the rigid system would have a lower Q than the flexible system). Above 10 Hz (the resonance for the flexibly-supported EDCs seem to be at 9.8 Hz) , we can see that the flexibly-supported EDC has slightly better isolation? I may need to take additional measurements of the transfer function of the flexibly-supported EDC (20 Hz to 100 Hz?)  to hopefully get a less-noisy transfer function at higher frequencies. The isolation does not appear to be that much better in the noisy region (above 20Hz). This may be because of the noise (possibly from the electromagnetic field from the shaker interfering with the magnets in the TT?). There is a 3rd resonance peak at about 22 Hz. I'm not sure what causes this peak...I want to confirm it with an FFT measurement of the flexibly-supported EDC (20 Hz to 40 Hz?)

 

 

EXPrigidvsflex.jpg

 

 

  5205   Fri Aug 12 11:07:50 2011 NicoleSummarySUSMore TT Shaking Completed This Morning

This morning (about 10am to 11am), I have collected additional transfer function measurements for the T.T. suspension. I have finished taking my measurements. The SR785 has been returned to its place next the the seismometer racks.

 The data has been backed up onto the cit40m computer

ELOG V3.1.3-